OpenStudy (anonymous):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution. x + y + z = -5 x - y + 3z = -1 4x + y + z = -2 A. {( 1, -4, -2)} B. {( -2, 1, -4)} C. {( 1, -2, -4)} D. {( -2, -4, 1)}

OpenStudy (anonymous):

first, make a matrix with the coefficients and constants ^_^

OpenStudy (anonymous):

i'll show you what the first row looks like to get you started 1 1 1 -5

OpenStudy (anonymous):

I don't really know how to do matrices. @DemolisionWolf

OpenStudy (anonymous):

so we are given these equations. x + y + z = -5 x - y + 3z = -1 4x + y + z = -2 lets look at the first one only x + y + z = -5 this can be written as this, with the coefficients of each term written in, (becuase when the coefficient is '1' we don't write it) x + y + z = -5 becomes 1x + 1y + 1z = -5 now, i make the first row of the matrix by taking the coefficients from each term and then the constant on the other side of the equal sign, which is 5 1 1 1 -5 the first '1' comes from the x, the next '1' comes from the y, and the third '1' comes from the z I'll show you quicker on the second equation given: x - y + 3z = -1 becomes 1x - 1y + 3z = -1 becomes 1 -1 3 -1 make sense?

OpenStudy (anonymous):

yes. so it should be: 1 1 1 -5 1 -1 3 -1 4 1 1 -2? @DemolisionWolf

OpenStudy (e.mccormick):

Yes, it would start with: \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 1 & -1 & 3 & -1\\ 4 & 1 & 1 & -2 \end{array}\right]\) I use ``` \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 1 & -1 & 3 & -1\\ 4 & 1 & 1 & -2 \end{array}\right]\) ``` to make it come out properly.

OpenStudy (anonymous):

Okay. So what's the next step?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

Know what upper echelon form or, row echelon form are? That is what you are going to want in the end. That is where the triangle has numbers and the lower part has zeros before the numbers. Let me just show it with letters because it is easier than trying to describe it. \(\left[\begin{array}{ccc|c} a & b & c & d \\ 0 & e & f & g\\ 0 & 0 & h & i \end{array}\right]\) To get it into that form, you do elementary row operations. Has your book or class covered those?

OpenStudy (anonymous):

Nope. They don't look familiar... @e.mccormick

OpenStudy (e.mccormick):

In short: Elementary Row Operations. 1. Interchange two rows. 2. Multiply a row with a nonzero number. 3. Add a row to another one. 4. Do 2 and 3 at the same time, so multiply a row by a number then add it to another one. So if you look at this: \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 1 & -1 & 3 & -1\\ 4 & 1 & 1 & -2 \end{array}\right]\) You want to get the 1 in the fist column of the second row to 0. What would happen if you multiplied the first row by -1, then added it to the second row to replace the second row?

OpenStudy (anonymous):

Multiply the first row by -1 meaning by 1,1,1, and -5?

OpenStudy (e.mccormick):

Yes. Each element in the row is multiplied by -1.

OpenStudy (anonymous):

So, I would get -1 -1 -1 and 5

OpenStudy (e.mccormick):

Good. And add that to the second row.

OpenStudy (anonymous):

Then, adding it to the 2nd row I'd get 0 -2 2 and 4

OpenStudy (e.mccormick):

=) Here I was working out the steps in case you did not get it, but you did! \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & -2 & 2 & 4\\ 4 & 1 & 1 & -2 \end{array}\right]\) OK, now, row 2 is done. Nothing else needs to be fixed for Gaussian Elimination. Row 3 needs both he first and 2nd columns changed to 0s. So, think you could do something similar?

OpenStudy (anonymous):

Multiply the 3rd row by a number.... but I'm not sure what number

OpenStudy (e.mccormick):

Actually, I would multiply the first row by a number and add it to the 3rd to get rid of the first column. Now, the first column is a 0. What would add to 4 to make 0?

OpenStudy (anonymous):

-4

OpenStudy (e.mccormick):

=) And there is what you need to multiply by.

OpenStudy (anonymous):

So multiply the first row by 4, then add to the 3rd row?

OpenStudy (anonymous):

-4*

OpenStudy (e.mccormick):

Yep!

OpenStudy (anonymous):

okay, so -4 -4 -4 and 20 when added to the third row, 0 3 3 and 18

OpenStudy (e.mccormick):

Watch your signs. =)

OpenStudy (anonymous):

Oh! 0 -3 -3 and 18

OpenStudy (e.mccormick):

OK: \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & -2 & 2 & 4\\ 0 & -3 & -3 & 18 \end{array}\right]\) Now, what could you multiply row 2 by to make the first non-zero number into a 1?

OpenStudy (anonymous):

1?

OpenStudy (e.mccormick):

Well, there is a -2 there...

OpenStudy (anonymous):

I thought by the first non-zero number you meant the 1

OpenStudy (e.mccormick):

In row 2.

OpenStudy (anonymous):

my mistake! I was looking at the rows.

OpenStudy (e.mccormick):

I forgot to mention something about the upper echelon form you use: It has 1 in the leading spots, so more like: \(\left[\begin{array}{ccc|c} 1 & b & c & d \\ 0 & 1 & f & g\\ 0 & 0 & 1 & i \end{array}\right]\) This is to get the 1 in the second row. you already have a 1 in the first row.

OpenStudy (anonymous):

Okay. So, I have to multiply the second row by a number that will make the -2 a 1

OpenStudy (e.mccormick):

Exactly!

OpenStudy (anonymous):

Can I multiply by a fraction?

OpenStudy (e.mccormick):

Yes. =) \(-\frac{1}{2}\) is valid, which I bet is what you wanted to use.

OpenStudy (anonymous):

yes! so that would then be 0 1 -1 and -2

OpenStudy (e.mccormick):

\(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & -1 & -2\\ 0 & -3 & -3 & 18 \end{array}\right]\) OK, so know what is next?

OpenStudy (e.mccormick):

To be honest, there is a choice of 2 things to do next. Either is fine.

OpenStudy (anonymous):

Could I add that to the third row?

OpenStudy (e.mccormick):

Add what to the 3rd row?

OpenStudy (anonymous):

the second row

OpenStudy (e.mccormick):

Well, would just the 2nd row get rid of it, or would you also need to multiply?

OpenStudy (anonymous):

I would alsp need to multiply

OpenStudy (anonymous):

also*

OpenStudy (e.mccormick):

OK! So get to it! Let me know what you get.

OpenStudy (anonymous):

okay. I would get 0 1 1 and -6

OpenStudy (e.mccormick):

Hmmm...

OpenStudy (e.mccormick):

\(3\left[\begin{array}{ccc|c} 0 & 1 & -1 & -2\\ \end{array}\right] + \left[\begin{array}{ccc|c} 0 & -3 & -3 & 18 \end{array}\right]\) becomes \(\left[\begin{array}{ccc|c} 0 & 3 & -3 & -6\\ \end{array}\right] + \left[\begin{array}{ccc|c} 0 & -3 & -3 & 18 \end{array}\right]\) When you add those, what do you get?

OpenStudy (anonymous):

0 0 -6 and 12

OpenStudy (e.mccormick):

=) \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & -1 & -2\\ 0 & 0 & 1 & -2 \end{array}\right]\) OK, that is in reduced upper echelon form. Now for back substitution. First, the last row is solved. \(\left[\begin{array}{ccc|c} 0 & 0 & 1 & -2 \end{array}\right]\) means \(y=-2\) So, this is where Gaussian Elimination gets a little odd. I'll show you with the second row, then you can try it on the first. OK?

OpenStudy (anonymous):

okay! =)

OpenStudy (e.mccormick):

Oops, I meant \(z=-2\) The 3rd column is z, the second y, and the first x. I put what is known as a multiplier into the matrix: \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & -1 & -2\\ 0 & 0 & 1 & -2 \end{array}\right]\) becomes \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & -1(-2) & -2\\ 0 & 0 & 1 & -2 \end{array}\right]\) Then I move it from the left to right using subtraction: \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & -1(-2)-[-1(-2)] & -2-[-1(-2)]\\ 0 & 0 & 1 & -2 \end{array}\right]\) becomes \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & 0 & -2-[2]\\ 0 & 0 & 1 & -2 \end{array}\right]\) becomes \(\left[\begin{array}{ccc|c} 1 & 1 & 1 & -5 \\ 0 & 1 & 0 & -4\\ 0 & 0 & 1 & -2 \end{array}\right]\) OK. So \(y=-4\). Can you now use both \(z=-2\) and \(y-4\) to do the same thing to x?

OpenStudy (anonymous):

yes, so that means I'd plug the z into the 2nd row and the y into the first row?

OpenStudy (e.mccormick):

You use the solved y and z to plug into the first row to start solving x.

OpenStudy (anonymous):

Okay. And by doing that would I get x=1?

OpenStudy (e.mccormick):

Yes!

OpenStudy (anonymous):

THANK YOUUU!!!!!!!!

OpenStudy (e.mccormick):

Want to see the whole thing as one post?

OpenStudy (anonymous):

No need! I've copied it all down =)

OpenStudy (e.mccormick):

OK. I put in some notation like: \(-1R_1+R_2 \rightarrow R_2\)

OpenStudy (e.mccormick):

It keeps it clear what was done.

OpenStudy (anonymous):

Thank you so much!!! Could I message you if I need help on anything else?

OpenStudy (e.mccormick):

Well, I am going to go home and get food. but you can tag me if you see me. If I don't respond, it means I am helping someone else or otherwise busy. Mods get tagged a lot. Hehe.

OpenStudy (anonymous):

haha okay no problem, thank you! Have a good night.

OpenStudy (e.mccormick):

You too. And if you see anyone with a purple A in front of their name, feel free to tag them too. They are my minion army.

OpenStudy (anonymous):

hahaha Thank you! Will do.